how would you graph
y=6/5x+8

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

I would pick 2 values of x and find y.
Then I'd plot those 2 points, and draw a line thru them.

Answer by mananth(16946)

y=6/5x+8
...
find the y & x intercepts
...
plug x=0
x=0,y=8 (0,8)
y=0,x=-20/3 (-20/3,0)
...

...
m.ananth@hotmail.ca